OpenStudy (scarlettthecamel):

Write 80% as a fraction in it's simplest form

OpenStudy (uri):

 Simplify it more @scarlettthecamel

OpenStudy (anonymous):

80% = 80/100 = 8/10 = 4/5 I believe that is correct.

OpenStudy (scarlettthecamel):

\[\frac{ 2 }{ 5 }\]

OpenStudy (scarlettthecamel):

oh

OpenStudy (anonymous):

4/5 would be in simplest form

OpenStudy (scarlettthecamel):

thx peeps xxx

OpenStudy (anonymous):

@LittleFirefly is correct

OpenStudy (anonymous):

No problem. :)

OpenStudy (anonymous):

your most certainly welcome :)